a rectangle with dimensions of 5 x 8 is inscribed in a circle. what is the area and circuference of this circle?[ inscribed means inside the vertexs touching the edge of the circle].
Answer by stanbon(75887)

A rectangle with dimensions of 5 x 8 is inscribed in a circle. What is the area and circumference of this circle?
Draw the picture.
Draw a diagonal of the rectangle.
The diagonal is also the diameter of the circle.
Notice the right triangle with the diagonal as the hypotenuse.
Use Pythagoras to find the length of the diagonal/diameter.
diag^2=5^2+8^2=89
diag=sqrt(89)=9.434
Circumference of the circle = (pi)d = 9.434(pi)=29.64 units
Area of the circle = (pi)r^2 = (pi)4.717^2=69.90 sq. units
Cheers,
Stan H.